Please provide detailed solution if possible
4d し、12A Determine: (No work need be shown) Rs 91。- lo V. ratio R、R,- 1:2
Answer #1

Consider the loop R1, R2, R6

given V1 = 20 and V2 = 60 , Vo = 90 supply

V1 + V2 + V6 = Vo =90

hence V6 = 10 V

V3 = V4 since R3 and R4 are parallel

V5 = 20 V given

V3 + V5 = V2 = 60 V

hence V3 = V4 = 60-20 = 40 V

Totel current entering (R3 + R4) = current leaving (R3+R4) = current through R5 = I5 = 6A

I4 = 2.0 A given

I4 +I3 = I5 = 6

I3 = 6-2 = 4.0A

R3/R4 = I4/I3 = 2/4 = 0.5

I6 = I2 +I5 ( sum of the current entering the node)

= 3.0 + 6.0 = 9.0 A

I1 = I2 + I3 + I6 ( total current entering the node is equal to that of leaving the node)

= 3.0 +4.0 +2.0 = 9.0 A